We are currently aware of a percentage of customers that are experiencing problems with ADSL connections, which may lead to disconnections or failure to route. Our engineers are working to resolve the issues as a matter of priority and we will update as soon as possible.
Archive for September, 2006
Our engineers are currently on site at our Wolverhampton POP and need to conduct an emergency reboot of the wolv.enta.net router to prevent failure. This work will take place at 11.30 this morning and will cause a small outage for directly connected customers. Anyone directly affected has been contacted and we foresee no problems for other customers. Please accept out apologies for the short notice and any inconvenience caused.
September 26, 2006 14:33
Our systems team are currently investigating reports of problems with some domains hosted on nbg.viper.enta.net.
Whilst this is ongoing we have disabled ftp access and request that no new domains or mailboxes are added to the viper. We will provide further updates as they are available and will endeavour to return services to normal as soon as possible.
Our systems team will this evening be undertaking some emergency work on the router bir.enta.net, increasing its available memory thereby requiring a reboot of the router. This will involve approximately 15 minutes downtime for all directly connected customers. We are currently attempting to contact all affected customers and apologise for any inconvenience the short notice may cause.
We are currently experiencing issues with our premier support line, resulting in calls not being delivered to our support staff. We request that all premier customers conact the standard support number if the need arises. The number to call is 0870 224 3494. We apologise for the problems and will advise once the premier number is available.
21st September 2006 22.00 – 23.00hrs
We will be performing a reboot of the redbus-sov3.core.enta.net router this evening between 22.00 and 23.00 to resolve a CPU load problem. Customers connected directly to this equipment will suffer a small period of downtime, somewhere in the region of 5 minutes. We are currently attempting to contact each customer that will be directly affected. We do not foresee any issues for other customers at this time.
Some ADSL users are currently experiencing a level of packet loss when connected to one of our 622 central pipes. This is being caused by a skewed distribution of sessions between the pipes, resulting in one pipe becoming congested, despite having sufficient capacity on the network. If you are suffering from packet loss, we suggest disconnecting and reconnecting your router to resolve the issue. We are working to ensure a fair distribution of sessions and will advise once the issue is resolved.
01st October 2006 08.00 – 16.00hrs
We have been advised by our mains electricity supplier that there will be a power outage across the Stafford Park area of Telford on 01st October 2006 between the hours of 08.00 – 16.00. This work is being completed to increase power supply resilience in the area. All services supplied from, or located in, our Telford HQ datacentres will be affected switching to UPS and backup power as required.
Our backup generators will automatically take over the generation and supply of power during the power outage. We have sufficient fuel to service this time period, and have also secured additional fuel on short notice should it be required for any reason. We will also be arranging a test of the automatic switchover before the date of the planned work and will advise according once the details have been finalised.
Thusday 21st September 2006 07.00 – 07.30hrs
We have been advised by LINX that they wish to move a 10Gb port utilised by Entanet from temporary to permanent equipment. The work is scheduled to be carried out at 07.00hrs on Thursday 21st September 2006 and should have little impact as network traffic will reroute around the area.
We are currently experiencing problems with routing in Telehouse East, which will be having knock on effects across the London ring. Our engineers are currently working to resolve this issue and may require a reboot of routing equipment in Telehouse East. We will endeavour to return services to normal as soon as possible.